OpenStudy (anonymous):

Your employer pays 85% of your health insurance premium. If your premium is $5236 annually, how much is your share of the premium?

OpenStudy (anonymous):

5236x.85= 4450.6 ( how much your employer pays) Simply subtract his 85% from the total of your premium to get your share

OpenStudy (anonymous):

Do proportions, so its asking your share so 100-85=15 \[\frac{ 15 }{ 100 }=\frac{ x }{ 5236 }\]Cross-multiply\[78540=100x\]Divide by 100 to get x by itself \[785.4=x\]

OpenStudy (anonymous):

5236-4450.6=785.4 ( your share)